Two small plastic spheres are given positive electrical charges. When they are a distance of 14.1 cm apart, the repulsive force between them has a magnitude of 0.205 N .

1. What is the charge on each sphere if the two charges are equal?

2. What is the charge on the first sphere if it equals one-quarter the charge of the second sphere?

3. What is the charge on the second sphere described in the previous part?
Answer #1

One Charge = q1 = q

Force = F = 0.205 N

Other charge = q2 =q

Distance = r = 14.1 cm = 0.141m

permittivity of free space = eo =8.854×10−12 C^2/ (N.m^2)

1.)

Using Coulomb's law,

F=[1/4pieo]q1q2/r^2

F = [1/4pieo]q^2 / r^2

q^2 =F [4pieo]r^2

q = sq rt [ 0.205[4pi8.85×10^−12 ] (0.141*0.141)

q =0.673*10^-6C


Each charge is 6.73*10^-7C ..Answer.
_______________________________________________

2.)

One Charge = q1 = q

Force = F = 0.205 N

Other charge = q2 =q1/4

Distance = r = 0.141 m

permittivity of free space = eo =8.854×10−12 C^2/ (N.m^2)

Using Coulomb's law,

F=[1/4pieo]q1q2/r^2

q1q2 =F [4pieo]r^2

(q)(q/4) = F[4pieo]r^2

q = sq rt { 0.205[4pi8.85×10^−12 ] (0.141*0.141)*4}

q = 1.35*10^-6 C


Therefore,charge on first sphere =q/4 =1.35*10^-6C =3.3710^-7 C ..Answer.

If the charge on first sphere is 1 fourth times the charge of the other.

_________________________________-

C)The charge on the second sphere is 1.35×10^-6 C ...Answer.

Hope this helps!.

Please thumbs up.